OpenStudy (anonymous):

5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is ___. -5 -1 1/2 1

OpenStudy (anonymous):

Do you not understand how to do these?

OpenStudy (anonymous):

Solve the system by substitution: y = 6 - 5x 5x + 3(6-5x) = -4 x = 22/10 Substitute back in the first equation, y = 6 - 5(22/10) = 6 - 11 = -5